An electron travels through the uniform magnetic field of field strength B = (2.5 i + 3.5 j ) mT and electric field of field strength 4.00 I V/M .   If the electron is moving with the velocity v= (1500 j + 2000 k)m/s. Calculate the net force acting on the electron in terms of unit vector notation.
Answer #1

B = 12.57 + 3,5 j) mt. E=41 V/M V = (15005 + 2002)mls. Elecbuie Foruce Fur CE = 1,6X16!ly 4 Fe = 6.4 x 10-19 11 N Magnetic